Consider the production function faced by an economy in a particular year, Y=square root of K. N. Assume that N equals to 1 what level of capital will be required to achieve output of 9million
1
Expert's answer
2019-02-01T11:00:07-0500

The production function is following:

Y=√K×N

where

Y-total output;

K – capital input.

If we know total output and N, we can find the required level of capital:

K=(Y/N)^2

K=(9,000,000/1)^2=81×〖10〗^12=81 trillion